OpenStudy (anonymous):

We have a box with 8 balls numbered: 0,0,0,1,1,1,2,2 we pick three balls together respectively, what's the prospect that the combination of the numbers of the balls we picked is 5 ?

OpenStudy (anonymous):

Knowing that the balls are shapely matched and we can't see them in the box

OpenStudy (welshfella):

how many ways can we get a sum of 5 with 3 balls?

OpenStudy (welshfella):

one way is 2,2 and one of the 1's right?

OpenStudy (welshfella):

there are 3 balls marked 1 so the total number of ways to get 5 is 3.

OpenStudy (anonymous):

There's actually only 1 way to get 5 because we pick up the ball, check out it's number and throw it away

OpenStudy (mathstudent55):

First, you need to find how many ways you can pick 3 balls that add up to 5. To have a sum of 5 with three balls, you must have the two balls labeled 2. That adds up to 4. Then you need a ball labeled 1. Since there are 3 balls labeled 1, there are 3 ways of picking 3 balls that add up to 5. Now you need to find how many ways there are of picking 3 balls out of 10 balls. For the first pick, you can choose from 8 balls. The second pick is out of 7 balls. The third pick is out of 6 balls. There are 8 * 7 * 6 = 336 ways of picking 3 balls. Of the 336 ways of picking 3 balls, 5 ways result in a sum of 5. P(sum of 5 with 3 balls) = 5/336

OpenStudy (anonymous):

The answer for P(sum of 5 with 3 balls) = 2/8 *1/7 *3/6 *3= 18/336 wanna know why that's the answer ?

OpenStudy (mathstudent55):

Yes, you are correct. You do need 2, 2, 1, but there are 18 ways of arranging those numbers. Since you need 2, 2, 1, the first 2 is 2 out of 8. The next 2 is 1 out of 7, then the 1 is 3 out of 6. Finally, once you have those picks, you can arrange them in 3 different ways. 2/8 * 1/7 * 3/6 * 3 = 18/336 = 3/56

OpenStudy (anonymous):

Yup correct

OpenStudy (mathstudent55):

Thanks for pointing it out.

OpenStudy (anonymous):

You're welcome :)